OpenStudy (anonymous):

simplify 1/2log[3](x)-2log[3](y)-4log[3](z)??? HELP PLEASE

OpenStudy (anonymous):

i think the answer is log[3] sqrt(xy^2)/z^4 but not for sure

OpenStudy (anonymous):

\[\LARGE \frac12\log_3(x)-2\log_3(y)-4\log_3(z)=\] \[\LARGE \log_3(x^{\frac12 })-\log_3(y^2)-\log_3(z^4)=\] \[\LARGE \log_3(\sqrt x)-\log_3(y^2)-\log_3(z^4)=\] now what would you do ? :)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

\[\LARGE log_a(b)-\log_a(c)=\log_{a}\left(\frac bc \right)\] just follow the rule, and you'll get the result...

OpenStudy (anonymous):

so its log[3](sqrtx/y^2z^4)?